OpenStudy (mathmath333):

What is the rate of simple interest for the first 4 year if the sum of rs 360 becomes rs 540 in 9 year and the rate of interest for the last 5 year is 6%? a. 4% b. 5% c. 3% d. 6%

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ 540=360+360\times \dfrac{4r+30}{100} }\end{align}\)